OpenStudy (anonymous):

z varies inversley with the product of x and y. If z=30 when x=3 and y=2, write the function that models the relationship.

OpenStudy (ybarrap):

z = k/(xy) 30=k/(3*2) k=30*12 Make sense?

OpenStudy (anonymous):

how did you get z = k/(xy) ?

OpenStudy (ybarrap):

Whenever something is proportional to something else, as in this case, there will be a proportionality constant that makes the problem exact. Here, we needed k because we know that z is inversely proportional to the product of x and y. This is like saying that z is proportional to 1/(x times y) or 1/(xy) or 1/xy. So now the question is, what is k? We are give values for z,x and y. Without the k, we can't get an exact answer.

OpenStudy (anonymous):

Thank you :)